# Causal Concept Graph Models: Beyond Causal Opacity in Deep Learning

### v4

@@ 5 Discussion
Related works
Causal CGMs present substantial advantages compared to the state of the art. Compared to most causal feature-attribution methods (e.g., (Chattopadhyay et al., 2019)), Causal CGMs focus on high-level human interpretable concepts. Causal CGMs differ from existing CBMs in their approach to intervention and causal relationships. The causal structure of the inferences that Causal CGMs deliver does not subsume weak forms of causal dependencies as in existing CBMs, where concepts are all direct causes of the target and causally independent of each other. This way, Causal CGMs enable human-in-the-loop corrections to mispredicted intermediate reasoning steps, boosting not just downstream accuracy after corrections but also the accuracy of the explanation provided for a specific instance. Closest to Causal CGMs in concept usage are post-hoc causal concept-based explainability techniques, like DiConStruct (de Oliveira Moreira et al., 2024) and conceptual counterfactual explanations (Abid et al., 2021). These methods build surrogate causal models to emulate a target black box model’s predictive behaviour. However, as Rudin (2019) notes, convergence in predictive behaviour does not ensure structural similarity in decision-making processes, making surrogate models as explanatory proxies questionable. Causally interpretable by-design architectures, such as Causal CGMs, do not suffer from this issue.
Limitations and future works
Our method’s limitations mainly stem from inherent constraints in CBMs and causal reasoning. The quality of learnt causal graphs depends on dataset quality and annotations, where missing or noisy labels can lead to suboptimal graphs. Similarly to generalised PGMs, Causal CGMs can be easily unfolded when the final graph is acyclic. Variables in cycles can still be inferred, but require special unfolding techniques which might be explored in future works. Finally, a Causal CGM’s learnt graph does not necessarily represent the causal mechanisms of the data-generating process, but rather that of the model’s inference. This makes our models suitable to be verified and controlled but not necessarily to understand their training data’s distribution.
@@ Appendix B Derivation of Causal CGMs
Definition B.1 (Generalized Causal Concept Graph Model).
Given an observed input feature $x$, a set of $k\in\mathbb{N}$ latent factors $u_{i}\in\mathcal{U}$ each associated with a high level interpretable variable $v_{i}\in\mathcal{V}$, a Generalized Causal Concept Graph Model is the generalised probabilistic graphical model (PGM) $G=(\mathcal{N},\mathcal{E})$ with nodes $\mathcal{N}=\{x\}\cup\mathcal{U}\cup\mathcal{V}$ and edges =$\mathcal{E}=\big{\{}(x,u_{i})\mid u_{i}\in\mathcal{U}\big{\}}\cup\big{\{}(u_{i},v_{i})\mid i\in\{1,\cdots,k\}\big{\}}\cup\{(v_{i},v_{j})\mid v_{i},v_{j}\in\mathcal{V},v_{i}\neq v_{j}\}$ which represents the joint conditional distribution $p(v,u\mid x)$:
@@ Appendix D Neural parametrization of Concept Graph Models
We can interpret the factors of a dissected Causal CGM as follows: $p(u_{j}\mid\mathbf{x};\theta_{\zeta})$ is the exogenous encoder, i.e., a deterministic distribution that is parametrised by a neural network $\zeta:X\to U$. In CBMs, this function represents the input encoder. The exogenous encoder $\zeta$ generates exogenous variables $u_{i}\in\mathcal{U}$ mapping raw input features $\mathbf{x}$ (e.g., an image’s pixels) to latent embeddings $\hat{\mathbf{u}}_{i}\in\mathbb{R}^{q},\ q\in\mathbb{N}$. In practice, this process mirrors the generation of context vectors in Concept Embedding Models to avoid information bottlenecks which may negatively affect the model’s accuracy (Espinosa Zarlenga et al., 2022). First, the encoder $\psi:X\rightarrow H$ maps raw features to a latent code $\mathbf{h}\in H$. Then, a pair of neural networks $\{\phi_{i}^{+},\phi_{i}^{-}\}$ map the latent code into two different embeddings whose concatenation $[\phi_{i}^{+}(\mathbf{h}),\phi_{i}^{-}(\mathbf{h})]^{T}$ corresponds to the exogenous variable $\mathcal{U}_{i}$ of the $i$-th concept:
@@
In order to learn explicit structural equations, existing logic-based aggregation methods can be used from the concept literature (Barbiero et al., 2022; 2023). App. F describes in more detail their adaptation in Causal CGM.
@@ Appendix F Architecture
Initialisation of adjacency matrix
Causal CGM provides versatile initialisation options for the adjacency matrix $A$, tailored to specific scenarios. In certain instances, weights can be derived from domain expertise or provided along with training data and labels. Without such information, weights can be directly inferred from training labels through causal structural learning algorithms (Kaddour et al., 2022) as a preliminary step. These approaches guide the model towards a predetermined decision-making pathway. Alternatively, weights can be learnt concurrently during the Causal CGM training phase, as outlined in Section 3.3. These weights may be initialised either randomly or based on the conditional entropy between labels, providing a better starting point. Additionally, a hybrid approach is feasible, where certain elements in $A$ are fixed while others remain trainable. For instance, a causal structural learning algorithm might yield a Partial Ancestral Graph (PAG) with undirected edges, allowing for the definition of directed edges and learning the direction for others to avoid cycle formation.
Causal mechanisms
In Causal CGM, the function $f_{i}$ corresponds to a causal mechanism in a SCM. Such mechanisms are typically formalised via structural equations. For instance, linear models are a common choice for label predictors in Concept Bottleneck Models (Koh et al., 2020) where endogenous embeddings are aggregated using a permutation invariant aggregator function $\oplus$ (such as the element-wise maximum, or sum):
@@
Compositional generalisation
The training procedure of Causal CGMs is highly parallelizable and modular as only direct connections need to be trained together (e.g., $a\rightarrow b$ and $b\rightarrow c$), while the model takes care of distant connections in an indirect way. For instance, the connection $a\rightarrow b\rightarrow c$ can be obtained as a composition of two different independent training procedures for $a\rightarrow b$ and $b\rightarrow c$. As a result, it is trivial for a Causal CGM to make the causal graph grow even at test time (see Figure 11). This can be done by composing two different graphs obtained by independent training procedures, encoders, datasets, or data types. We note that this is not possible in standard CBMs, which need to re-train the task predictor from scratch whenever new concepts or tasks are added to the mix. Moreover, this modularity enables a form of out-of-distribution compositional generalisation as it creates new distant connections between variables that were never part of the same training procedure (e.g., $a$ and $c$ in the previous example).
@@ G.4 Implementation details
Hyperparameters All baseline and proposed models were trained for varying epochs across different datasets: 500 for Checkmark, 200 for dSprites, 30 for CelebA and 25 CIFAR10. The optimal epoch for each was determined based on label accuracy on the validation set. A uniform learning rate of 0.01 was applied across all models and datasets. For the CBM and CEM models, both concept and task losses were equally weighted at 1. This weighting scheme was also applied to the loss terms for endogenous copies’ prediction, endogenous variables’ prediction ($\lambda_{1}$), and graph priors ($\lambda_{2}$). The weight assigned to the loss terms in our models to maximize CaCE is 0.05. Additionally, $\gamma$ was treated as a learnable parameter, initialized at 0.1, and $\beta$ was set to 1. All experiments were conducted using five different seeds (1, 2, 3, 4, 5).
Code, licenses and hardware For our experiments, we implement all baselines and methods in Python 3.9 and relied upon open-source libraries such as PyTorch 2.0 (Paszke et al., 2019) (BSD license), PytorchLightning v2.1.2 (Apache Licence 2.0), Sklearn 1.2 (Pedregosa et al., 2011) (BSD license). In addition, we used Matplotlib (Hunter, 2007) 3.7 (BSD license) to produce the plots shown in this paper. Two datasets we used are freely available on the web with licenses: dSprites (Apache 2.0) and CelebA, which is released for non-commercial use research purposes only. We also introduce the Checkmark dataset and we described it in this section. We will publicly release the code with all the details used to reproduce all the experiments under an MIT license. All the experiments except the CIFAR10 ones were performed on a device equipped with an M3 Max and 36GB of RAM, without the use of a GPU. The CIFAR10 experimentswere conducted on a workstation equipped with an NVIDIA RTX A6000 GPU, two AMD EPYC 7513 32-Core processors, and 512 GB of RAM.Approximately 60 hours of computational time were utilized from the start of the project, whereas reproducing the experiments detailed here requires only 8 hours.
@@ H.6 $k$-unfolding ablation
In Causal CBM, $1$-step unfolding of the causal graph during training is equivalent to any k-step ($k>1$) unfolding as any step corresponds (theoretically and empirically) to repeating the same query on the same conditional probability table (CPT). Causal CGM learns the conditional probabilities of the variables given its parents (i.e. the CPT of the Bayesian network). Under full observability, conditional queries of the form $p(v|\text{ all but }v)$ are just inspections of the CPT and do not require any reasoning on the cyclical model (which alleviates us from setting a complex $k$-step semantics to unroll the cyclical model). The key observation is that the cyclical model served the purpose of templating multiple (any) acyclical graphs jointly without making assumptions on concept relations. During learning, the probabilities of the CPT (parameterized by NNs) are also subject to an acyclicity objective. After learning, the resulting graph is acyclic thus inference is performed via standard message passing along the directed acyclic graph. To show this experimentally, we report the results on our CBM benchmark using a different number of unfolding steps and considering different levels of concept completeness in Table 10. As expected, the results show that increasing the number of steps does not impact Causal CGM’s classification performances.
@@ Appendix I Answering counterfactual queries with Causal Concept Graph Models
Causal CGMs are not identifiable in the sense that we can obtain different causal graphs by training the model over the same dataset. This is similar to obtaining different parameter values and embeddings when training standard deep-learning black boxes on the same data. Similarly, the training of causal CGMs can result in different causal decision-making structures for the same task. The difference is that in the case of the black box, we cannot compare the different models because they are (causally) opaque. In the case of Causal CGMs, we can inspect the multiple models obtained and verify them against desirable properties (e.g., fairness), hence adopting the best one.

### v6

@@ 5 Discussion
Related works Causal CGMs present substantial advantages compared to the state of the art. Compared to most causal feature-attribution methods (e.g., (Chattopadhyay et al., 2019)), Causal CGMs focus on high-level human-interpretable concepts. Causal CGMs differ from existing CBMs in their approach to intervention and causal relationships. Unlike vanilla CBMs, where concepts are direct, independent causes of the target, Causal CGMs incorporate richer causal structures. This enables human-in-the-loop corrections to mispredicted reasoning steps, improving both downstream accuracy and explanation fidelity. Causal CGMs are also more flexible than existing relational CBMs (Barbiero et al., 2024) where concept relations are assumed to be provided a priori. The closest methods to Causal CGMs are post-hoc causal concept-based explainability techniques, like DiConStruct (de Oliveira Moreira et al., 2024) and conceptual counterfactual explanations (Abid et al., 2021). These use surrogate causal models to emulate a black box’s predictions. However, as Rudin (2019) notes, matching predictive behaviour does not guarantee structural similarity in decision-making, making surrogate models unreliable explanatory proxies. Causally interpretable by-design architectures, such as Causal CGMs, do not suffer from this issue.
Limitations and future works Our method’s limitations stem from those of CBMs and causal reasoning. The quality of learnt causal graphs depends on dataset quality and annotations; missing or noisy labels may lead to suboptimal graphs. Like generalised PGMs, Causal CGMs unfold easily when the final graph is acyclic. Cyclic variables remain inferable but require special unfolding techniques, which future work may explore. Finally, a Causal CGM’s learnt graph represents the model’s inference, not necessarily the data-generating process, making it suitable for verification and control but not for understanding the dataset’s distribution.
@@ 6 Conclusion
Acknowledgments
GD acknowledges support from the European Union’s Horizon Europe project SmartCHANGE (No. 101080965). PB acknowledges support from Swiss National Science Foundation projects TRUST-ME (No. 205121L_214991) and IMAGINE (No. TMPFP2_224226). MEZ acknowledges support from the Gates Cambridge Trust via a Gates Cambridge Scholarship. AT acknowledges the support by the Hasler Foundation grant Malescamo (No. 22050), and the Horizon Europe grant Automotif (No. 101147693). MG acknowledges support from Swiss National Science Foundation projects XAI-PAC (No. PZ00P2_216405). GM acknowledges support from the KU Leuven Research Fund (STG/22/021, CELSA/24/008) and from the Flemish Government under the "Onderzoeksprogramma Artificiële Intelligentie (AI) Vlaanderen" programme.
@@ Appendix B Derivation of Causal CGMs
Definition B.1 (Generalised Causal Concept Graph Model).
Given an observed input feature $x$, a set of $k\in\mathbb{N}$ latent factors $u_{i}\in\mathcal{U}$ each associated with a high level interpretable variable $v_{i}\in\mathcal{V}$, a Generalised Causal Concept Graph Model is the generalised probabilistic graphical model (PGM) $G=(\mathcal{N},\mathcal{E})$ with nodes $\mathcal{N}=\{x\}\cup\mathcal{U}\cup\mathcal{V}$ and edges =$\mathcal{E}=\big{\{}(x,u_{i})\mid u_{i}\in\mathcal{U}\big{\}}\cup\big{\{}(u_{i},v_{i})\mid i\in\{1,\cdots,k\}\big{\}}\cup\{(v_{i},v_{j})\mid v_{i},v_{j}\in\mathcal{V},v_{i}\neq v_{j}\}$ which represents the joint conditional distribution $p(v,u\mid x)$:
@@ Appendix D Neural parametrization of Concept Graph Models
We can interpret the factors of a dissected Causal CGM as follows: $p(u_{j}\mid\mathbf{x};\theta_{\zeta})$ is the exogenous encoder, i.e., a deterministic distribution that is parametrised by a neural network $\zeta:X\to U$. In CBMs, this function represents the input encoder. In causal CGMs we use the exogenous encoder to predict the value of exogenous variables as we consider that the observed input (e.g., an image’s pixels) holds most of the contextual variability required to infer exogenous variables. For instance, an image provides information about background, lighting conditions, and object locations. All this information is used to anchor the endogenous predictor to a specific context and to correctly infer the values of the exogenous variables and the endogenous roots of the causal graph. The exogenous encoder $\zeta$ generates exogenous variables $u_{i}\in\mathcal{U}$ mapping raw input features $\mathbf{x}$ (e.g., an image’s pixels) to latent embeddings $\hat{\mathbf{u}}_{i}\in\mathbb{R}^{q},\ q\in\mathbb{N}$. In practice, this process mirrors the generation of context vectors in Concept Embedding Models to avoid information bottlenecks which may negatively affect the model’s accuracy (Espinosa Zarlenga et al., 2022). First, the encoder $\psi:X\rightarrow H$ maps raw features to a latent code $\mathbf{h}\in H$. Then, a pair of neural networks $\{\phi_{i}^{+},\phi_{i}^{-}\}$ map the latent code into two different embeddings whose concatenation $[\phi_{i}^{+}(\mathbf{h}),\phi_{i}^{-}(\mathbf{h})]^{T}$ corresponds to the exogenous variable $\mathcal{U}_{i}$ of the $i$-th concept:
@@
In order to learn explicit structural equations, existing logic-based aggregation methods can be used from the concept literature (Barbiero et al., 2022; 2023; Debot et al., 2024). App. F describes in more detail their adaptation in Causal CGM.
@@ Appendix E Unfolding CGMs
In rare cases, some edges with very small weights may still make the adjacency matrix $A$ cyclic. At inference time, we address this by iteratively removing the edges with the smallest weights from the cycles until the graph becomes a DAG.
@@ Appendix F Architecture
Initialisation of adjacency matrix Causal CGM provides versatile initialisation options for the adjacency matrix $A$, tailored to specific scenarios. In certain instances, weights can be derived from domain expertise or provided along with training data and labels. Without such information, weights can be directly inferred from training labels through causal structural learning algorithms (Kaddour et al., 2022) as a preliminary step. These approaches guide the model towards a predetermined decision-making pathway. Alternatively, weights can be learnt concurrently during the Causal CGM training phase, as outlined in Section 3.3. These weights may be initialised either randomly or based on the conditional entropy between labels, providing a better starting point. Additionally, a hybrid approach is feasible, where certain elements in $A$ are fixed while others remain trainable. For instance, a causal structural learning algorithm might yield a Partial Ancestral Graph (PAG) with undirected edges, allowing for the definition of directed edges and learning the direction for others to avoid cycle formation.
Causal mechanisms In Causal CGM, the function $f_{i}$ corresponds to a causal mechanism in a SCM. Such mechanisms are typically formalised via structural equations. For instance, linear models are a common choice for label predictors in Concept Bottleneck Models (Koh et al., 2020) where endogenous embeddings are aggregated using a permutation invariant aggregator function $\oplus$ (such as the element-wise maximum, or sum):
@@
Compositional generalisation The training procedure of Causal CGMs is highly parallelizable and modular as only direct connections need to be trained together (e.g., $a\rightarrow b$ and $b\rightarrow c$), while the model takes care of distant connections in an indirect way. For instance, the connection $a\rightarrow b\rightarrow c$ can be obtained as a composition of two different independent training procedures for $a\rightarrow b$ and $b\rightarrow c$. As a result, it is trivial for a Causal CGM to make the causal graph grow even at test time (see Figure 11). This can be done by composing two different graphs obtained by independent training procedures, encoders, datasets, or data types. We note that this is not possible in standard CBMs, which need to re-train the task predictor from scratch whenever new concepts or tasks are added to the mix. Moreover, this modularity enables a form of out-of-distribution compositional generalisation as it creates new distant connections between variables that were never part of the same training procedure (e.g., $a$ and $c$ in the previous example).
@@ G.2 Baselines
In our experiments, all baselines are trained using a joint training technique commonly used in CBMs (Koh et al., 2020). In joint training, the model is trained end-to-end with the task predictor directly taking the concept encoder’s outputs as input, enabling simultaneous optimization of both concept and task predictions.
@@ G.4 Implementation details
Hyperparameters All baseline and proposed models were trained for varying epochs across different datasets: 500 for Checkmark, 200 for dSprites, 30 for CelebA and 25 CIFAR10. The optimal epoch for each was determined based on label accuracy on the validation set. A uniform learning rate of 0.01 was applied across all models and datasets. For the CBM and CEM models, both concept and task losses were equally weighted at 1. This weighting scheme was also applied to the loss terms for endogenous copies’ prediction, endogenous variables’ prediction ($\lambda_{1}$), and graph priors ($\lambda_{2}$). The weight assigned to the loss terms in our models to maximise CaCE is 0.05. Additionally, $\gamma$ was treated as a learnable parameter, initialised at 0.1, and $\beta$ was set to 1. All experiments were conducted using five different seeds (1, 2, 3, 4, 5).
Code, licenses and hardware For our experiments, we implement all baselines and methods in Python 3.9 and relied upon open-source libraries such as PyTorch 2.0 (Paszke et al., 2019) (BSD license), PytorchLightning v2.1.2 (Apache Licence 2.0), Sklearn 1.2 (Pedregosa et al., 2011) (BSD license). In addition, we used Matplotlib (Hunter, 2007) 3.7 (BSD license) to produce the plots shown in this paper. Two datasets we used are freely available on the web with licenses: dSprites (Apache 2.0) and CelebA, which is released for non-commercial use research purposes only. We also introduce the Checkmark dataset and we described it in this section. We will publicly release the code with all the details used to reproduce all the experiments under an MIT license. All the experiments except the CIFAR10 ones were performed on a device equipped with an M3 Max and 36GB of RAM, without the use of a GPU. The CIFAR10 experimentswere conducted on a workstation equipped with an NVIDIA RTX A6000 GPU, two AMD EPYC 7513 32-Core processors, and 512 GB of RAM.Approximately 60 hours of computational time were utilised from the start of the project, whereas reproducing the experiments detailed here requires only 8 hours.
@@ H.6 $k$-unfolding ablation
In Causal CBM, $1$-step unfolding of the causal graph during training is equivalent to any k-step ($k>1$) unfolding as any step corresponds (theoretically and empirically) to repeating the same query on the same conditional probability table (CPT). Causal CGM learns the conditional probabilities of the variables given its parents (i.e. the CPT of the Bayesian network). Under full observability, conditional queries of the form $p(v|\text{ all but }v)$ are just inspections of the CPT and do not require any reasoning on the cyclical model (which alleviates us from setting a complex $k$-step semantics to unroll the cyclical model). The key observation is that the cyclical model served the purpose of templating multiple (any) acyclical graphs jointly without making assumptions on concept relations. During learning, the probabilities of the CPT (parameterised by NNs) are also subject to an acyclicity objective. After learning, the resulting graph is acyclic thus inference is performed via standard message passing along the directed acyclic graph. To show this experimentally, we report the results on our CBM benchmark using a different number of unfolding steps and considering different levels of concept completeness in Table 10. As expected, the results show that increasing the number of steps does not impact Causal CGM’s classification performances.
@@
H.7 Debiasing via do-interventions
We can measure the effectiveness of blocking in debiasing Causal CGMs using the following procedure:
  • Selection of Variables $i$ and $j$: We randomly select a pair of concepts, $i$ (parent) and $j$ (child), connected in the graph discovered by the CD algorithm or CausalCGM.
  • Do-Intervention: A do-intervention is applied to the child node ($j$) to cut off its dependence on the parent node ($i$). In our model, this intervention ensures that all subsequent nodes in the graph become independent of the parent node, as information flow is blocked by design.
  • Measuring the Impact: After the intervention, we modify the value of the parent node ($i$) and measure its impact on the final label (leaf node) with $i$ as an ancestor. If no effect is observed, it indicates successful debiasing, as expected by our model’s design. Cutting the edge breaks the information flow through the graph—from root to leaf—preventing the parent node from influencing subsequent nodes.
  • Comparison with Baselines: Baselines have no graph constraints, therefore the do-intervention modifies the value of a concept but does not alter the decision-making process. As a result, the parent node’s influence cannot be entirely removed. This contrasts our model’s ability to block the information flow fully.
The results of this experiment are reported in Table 3 using the Residual Concept Causal Effect ($\downarrow$) between causally-related variables having blocked all paths between the two variables with do-interventions on the causal graph. The optimal value is zero corresponding to perfect causal independence. To explain why the optimal value is zero, consider the following example. Consider a graph $A\rightarrow B\rightarrow C$. For this graph, the variable $A$ is a cause of the variable $C$. As a result, we expect the average causal effect of $A$ on $C$ to be higher than 00. Our experiment aims to show that intervening on $B$ makes $A$ and $C$ causally independent (cf. with L323-331). To show this, we apply a do-intervention in $B$, removing the edge $A\rightarrow B$ and generating two disconnected graphs: $A$ and $B\rightarrow C$. Now, we expect that the residual average causal effect of $A$ on $C$ is 00 because there is no longer a connection between $A$ and $C$. As a result, 00 is the optimal desirable outcome.
@@ Appendix I Answering counterfactual queries with Causal Concept Graph Models
Causal CGMs are not identifiable in the sense that we can obtain different causal graphs by training the model over the same dataset. This is similar to obtaining different parameter values and embeddings when training standard deep-learning black boxes on the same data. Similarly, the training of causal CGMs can result in different causal decision-making structures for the same task. The difference is that in the case of the black box, we cannot compare the different models because they are (causally) opaque. In the case of Causal CGMs, we can inspect the multiple models obtained and verify them against desirable properties (e.g., fairness), hence adopting the best one. This distinction impacts key properties, including identifiability. To further clarify this point, consider the following example. In the world, there is a single ground truth causal structure (e.g., “smoking causes lung cancer”). Guaranteeing identifiability is crucial here because it ensures that we can reliably recover the ground truth causal graph. In contrast, causal opacity deals with DL models, which are systems that can implement a variety of decision-making mechanisms (e.g., a valid decision mechanism for a DL model is: “the model predicts that an individual has lung cancer and this causes the model to predict that the individual has a smoking habit”). These mechanisms may differ even when trained on the same data, as DL models can learn different representations to predict the same outputs. As a result, there is no single "ground truth" causal structure to identify—multiple valid causal graphs exist, each reflecting a different way the model processes information. This inherent variability makes identifiability irrelevant in the context of causal opacity.